# How important is the dielectric constant in water modeling? Evaluation of the performance of the TIP4P/ force field and its compatibility with the Joung-Cheetham NaCl model

### v1

@@ Abstract
Efficient large scale computer simulations of complex aqueous solutions require the use of accurate but simple empirical force fields for water. However, the growing complexity of these systems is evidencing the difficulties to describe solution properties without due account of polarization. Conflicting strategies to remedy this problem are parametrizing water force fields to the dielectric constant or charge scaling of the solvated ions. In this work we compare results from TIP4P/ and OPC models, which are parametrized with an emphasis in the prediction of the dielectric constant, with results from TIP4P/2005, which is closer in spirit to the charge scaling strategy. The overall performance of the models is rated according to the Vega-Abascal benchmark. Our results show that TIP4P and TIP4P/2005 perform equally good, with the OPC model lying significantly behind. TIP4P/ can predict bulk phase properties (transport properties, thermal expansion coefficients, densities) of both liquid water and ice polymorphs, as well as surface tensions, with an accuracy very similar to TIP4P/2005, while performing very well for dielectric constants over a wide range of pressures and temperatures. On the other hand, TIP4P/2005 provides a better description of phase boundaries, including liquid-vapor and freezing transitions. However, the accurate prediction of dielectric constants allows TIP4P/ to describe solution densities for model ions targeted to their crystal and melt properties. This is achieved without the need to rescale charges, modify the Lorentz-Berthelot rule or tune the ion’s Lennard-Jones parameters. Our findings hinge on the significance of dielectric constants as a target property and cast some doubt on the need for the appealing concept of charge scaling.
keywords:
electrolyte solutionsDepartamento de Química Física, Facultad de Ciencias Químicas, Universidad Complutense, Madrid, 28040, Spain.
1 Introduction
@@
Water is essential for the existence of life on Earth. A simple molecule, yet exhibiting numerous anomalies, such as a negative expansion coefficient below 4 ∘C, a sharp increase of response functions upon cooling, a decrease of viscosity with the increase in pressure, and many others 1.
@@
This unique behavior makes water a perfect testbed for the study of a wide range of fundamental physical principles in statistical mechanics. Therefore, great efforts have been devoted to the development of accurate and meaningful force fields ever since the advent of computer simulations 2.
@@
On physical grounds, it is clear that the design of an all-purpose model which can adapt depending on the environment requires dealing with the polarizability issue, if not exactly, at least approximately. Unfortunately, polarizable models such as the BK3 3, HBP 4, iAMOEBA 5 give only a modest improvement, but their computational cost makes them inefficient for large system sizes. On the other hand, the very active line of research based on machine learned potentials can do no better than the underlying approximations 6, and some recent studies for state of the art density functionals provide very disappointing results 7, 8. Further improvement along this line requires a training set obtained from expensive ab-initio calculations dealing accurately with electronic correlations, but alas, neural networks appear to exhibit here significant limitations to model the ab-initio potential energy surface for a wide range of different chemical environments 9.
@@
Due to these difficulties, model development is still often directed to the design of simple and tractable force fields, based on intuition and basic principles of chemistry. As a result, one often neglects polarizability altogether, and describes the average molecular polarity with a simple distribution of fixed point charges embedded into a spherical core. This is the case of the celebrated SPC model 10. Here, two positive partial charges sitting on hydrogen positions are neutralized by a negative charge located at the center of a simple Lennard-Jones site, which accounts for repulsion and dispersion of the oxygen atom. Alternatively, the negative charge can be shifted along the bisector axis of the HOH angle, as proposed by Bernal-Fowler already in 1933 11, leading to the so-called TIP4P family of water models.12
@@
Despite their simplicity, SPC and TIP4P models do a rather good job at describing the qualitative behavior of liquid water at ambient conditions, but the TIP4P family has been shown to be more robust, and remains qualitatively correct over a much wider range of thermodynamic conditions 13, 14. Particularly, the TIP4P/2005 model provides a good qualitative description of the phase diagram, including the coexistence between several solid phases, as well as a quantitative description of the temperature of maximum density, the orthobaric liquid densities, molar volumes and surface tensions 15, 14.
@@
Of course, neglecting polarizability comes at a price. For the particular case of the TIP4P/2005 model, the accuracy of liquid phase densities is at the cost of a rather poor description of the dielectric constant, which is about 30% too low compared to experiments 16. Alternatively, it often occurs that models providing a good description of the dielectric constant, such as SPC/E or TIP3P 10, 12, yield a qualitatively incorrect description of the phase diagram 16.
@@
A simple way out of this problem was suggested by Leontyev and Stuchebrukhov some time ago 17, 18. These authors argued that the electric field generated by point charges on a molecule is screened by the electronic polarizability of the surrounding molecules. Therefore, the ”effective” charges, that are required to reproduce an accurate force field are smaller than the ”true” point charges of the molecule, by a factor equal to the high frequency refractive index, , i.e. . Consequently, models predicting a good force field will yield too low a molecular dipole moment when measured in terms of the effective charge. However, assuming the true charge is a factor larger than the effective charge, and noting water’s refractive index is ca. , brings the dipole moment of usual force fields ( D) in close agreement with results of ab-initio calculations ( D).
@@
This appealing idea has received great attention and has become known as the scaled charge paradigm 19, 20, 21, 22, 23, 24, 25. As another side of the same coin, some authors have stressed the need to make a distinction between the potential energy surface, as governed by , and the so called dipole moment surface, which is dictated by 26. This concept can be useful to bring predicted dielectric constants into agreement with experiment 27, 23, but must be exercised with great caution. Indeed, these two surfaces could be different but are definitively not independent. Linear response theory implies that a molecule cannot tell the difference between an external electric field (as dictated by a test charge) or that emanating from a neighboring molecule within the system (as dictated by the effective charge). However, in the scaled charge paradigm the dipolar response would be dictated by the true dipole in the former case and by the scaled dipole in the latter. As a result, arbitrarily scaling the molecular dipole to bring the dielectric constant into agreement with experiments is likely to violate the dissipation-fluctuation theorem.
@@
Be as it may, the charge scaling concept serves to increase the number of liquid state properties that can be predicted by non-polarizable force fields, and optimized models such at TIP4P/2005 can be considered as very robust tools for the prediction of pure water properties. However, the lack of polarization will result in poor performance in cases where environments with widely differing structure are encountered in the same system 14.
@@
One such important situation is when two or more phases need to be considered simultaneously, as in phase coexistence. Indeed, it has been widely documented that it is not possible to predict accurate liquid water properties while providing a good estimate of the melting point 14. Similarly, point charge models provide usually rather poor vapor pressures, even when including ad hoc corrections for the free energy of polarization 10, 28.
@@
A second very significant case involves electrolyte solutions, where ions can create very strong local electric fields and polarization becomes then a major issue. Indeed, for a single isolated ion of charge , the free energy of solvation is proportional to 29, so one can anticipate that the lack of an accurate dielectric constant will result in a poor description of the solution. Not surprisingly, attempts to reproduce electrolyte solutions using the expected nominal charge have met only limited success 30, 31, 32, 33.
@@
The scaled charge concept comes here at rescue, under the assumption that also the nominal charge of the ions needs to be scaled 19. Indeed, several authors have recently exploited this idea 19, 20, 21, 22, 23, 34, and showed that some of the TIP4P models with poor predictions of the dielectric constant can provide rather accurate predictions of a number of solution properties by scaling the nominal charges by a factor of ca. 22, 35, 36, 25. Unfortunately, predictions here are far less robust than for pure water, and different scaling factors need to be considered depending on whether one seeks for accurate predictions of bulk thermodynamic properties, surface tensions or transport properties 25, 37. Moreover, the improvement of the solution properties comes at the cost of significantly deteriorating the description of the melt and the crystal, which also results in poor predictions of the electrolyte’s solubility.
@@
For this reason, some authors have sought for an improved description of solution properties by using water’s dielectric constant as a major target property in the model parametrization 38, 39
@@
This is the case of the OPC model38, which has been recently recommended for use with the Amber package 40. Unfortunately, the improved description of the dielectric constant in this model is at the cost of using the location of the hydrogen sites as fitting parameters. This can provide extra flexibility for the parametrization, but will obviously ruin the molecule’s moment of inertia, which is one of the very few molecular properties that we can safely constrain from the outset. Of course, the dynamics will suffer from such a choice, and recent work has shown that the increased parameter space chosen is not sufficient to improve other models such as TIP4P/2005 41. On the other hand, TIP4P/, a model developed by Fuentes-Azcatl et al. at about the same time 39, reproduces experimental dielectric constants over a wide range of thermodynamic conditions, with only small variations of the TIP4P/2005 charges and geometry of the parent model. As a result, it can also predict accurately the location of the temperature of maximum density. The simultaneous prediction of these two properties is rather uncommon in point charge force fields, and appears to suggest a good performance of solution properties.
@@
Based on the scaled charge paradigm, however, it would appear that using dielectric constants as a target property is unsafe and will result in a poor description of the”true” force field 41. According to this argument, one would then expect that the improved performance will be at the cost of spoiling other properties, as it appears to be the case for the OPC model 41.
@@
In this work, we assess the performance of the TIP4P/ water model for a wide range of different properties, and compare its performance with OPC and TIP4P/2005 using the well known benchmark proposed by Abascal and Vega (VA) 16. The model is then tested for transferability by studying solution properties as described by the Joung-Cheetham electrolyte force field 30, which was not specifically parametrized for TIP4P/. Our results show that TIP4P/ performs essentially as well as TIP4P/2005 for pure water, and that it remains a transferable and robust model for the study of electrolyte solutions without additional parametrization. This casts some doubts on the otherwise appealing concept of charge scaling.
2 Overview of charge scaling
@@
Such a continuum description makes sense provided the distance between the interacting charges is large compared with the typical distances between the polarizable solvent molecules. Charge scaling is therefore the exact long wavelength result for the interaction between ions in a solvent. Assuming that interactions between nearby ions or solvent molecules is mediated by a dielectric continuum is likely to be significantly less accurate, however.
@@
Assuming an isotropic, local and linear response of the dielectric background, the polarization is , where is the background’s susceptibility. In as much as the nuclei are fixed, and the polarization is only due to the electronic degrees, we can assume is dictated by the static electronic polarization of the system. In practice, electronic degrees of freedom are unresponsive below optical frequencies 42, 29. Therefore, the response of the medium is dictated by the system’s index of refraction in the visible, . This is convenient, because the nuclei move in the scale of Infra-Red frequencies at most, so that the electrostatic approximation holds to a good approximation.
@@
This is the rationale of charge scaling. Since the pair energy is dictated by the scaled charges, , and the dielectric constant measures fluctuations of the actual charges, , it would appear to be unwise to use dielectric constants as a target property in force field modeling.
@@
with the non-local electronic response given by .
Borrowing ideas from statistical mechanics and field theory 43, 44, 45, 46, 47, we expect that a minimal model for the the wave-vector dependent refractive index could be given by:
where is a correlation length of the order of the solvent molecular spacing.
Fourier transforming Eq.(4), using Eq.(5) for the polarization response yields an effective pair potential for the electronic background mediated interactions (c.f. 48):
This result smoothly interpolates between Coulomb’s law in vacuum and a screened Coulomb’s law for charges in a polarizable background. The crossover takes place at a microscopic length scale of . This result might be a mere caricature of the complex distance dependent screening in water 44, 45, 49, but helps showing how the simplest account of non-locality can lead to deviations from the continuum approximation. Interestingly, Eq.(6) can be also interpreted as describing a plain Coulomb law with an dependent permitivity of , which smoothly switches from unscreened interactions to the expected long wave-length limit for large . The crucial issue here is what is the extent of charge screening at the typical distance of closest approach between the molecules, say, . i.e., whether is already sufficiently small for full screening to have set on. Since we expect on the scale of , this is not likely to be the general situation. This is in line with recent ab-initio calculations between ions dissolved in argon, which show that the screening at contact distances is not quite that achieved at large distances, but at least, definitively much larger than that expected for ions in vacuum 50. This means that charge screening is not fully warranted at contact distances, but could be a better approximation than assuming charges interacting in vacuum.
3 Methods
3.1 Dielectric constant

### v2

# How important is the dielectric constant in water modeling? Evaluation of the performance of the TIP4P/ force field and its compatibility with the Joung-Cheatham NaCl model

@@ Abstract
Efficient large-scale computer simulations of aqueous solutions require the use of accurate but simple empirical force fields for water. However, the complexity of these systems evidences the difficulties in describing solution properties without due account of polarization. Different strategies to remedy this problem are parametrizing water force fields to the dielectric constant or charge scaling of solvated ions. In this work, we compare results from TIP4P/ and OPC models, which are parametrized to predict the dielectric constant, with results from TIP4P/2005, which is closer in spirit to the charge scaling strategy. The performance of the models is rated according to the Vega-Abascal benchmark. Our results show that TIP4P/ and TIP4P/2005 perform equally well, with the OPC model lying significantly behind. TIP4P/ can predict bulk phase properties (transport properties, thermal expansion coefficients, densities) of both liquid water and ice polymorphs, but also surface tensions, with an accuracy very similar to TIP4P/2005, while performing very well for dielectric constants over a wide range of pressures and temperatures. On the other hand, TIP4P/2005 provides a better description of phase boundaries, including liquid-vapor and freezing transitions. However, the accurate prediction of dielectric constants allows TIP4P/ to describe densities of NaCl solutions for models parametrized to their crystal and melt properties only. This is achieved without the need to rescale charges, modify the Lorentz-Berthelot rule or tune the ion’s Lennard-Jones parameters. Our findings hinge on the significance of dielectric constants as a target property and show that a robust parametrization can be achieved without invoking the concept of charge scaling.
I Introduction
@@
Water is essential for the existence of life on Earth. A simple molecule, yet exhibiting numerous anomalies, such as a negative expansion coefficient below 4 ∘C, a sharp increase of response functions upon cooling, a decrease of viscosity with the increase in pressure, and many others Gallo et al. (2016).
@@
This unique behavior makes water a perfect testbed for the study of a wide range of fundamental physical principles in statistical mechanics. Therefore, great efforts have been devoted to the development of accurate and meaningful force fields ever since the advent of computer simulations Ouyang and Bettes (2015).
@@
On physical grounds, it is clear that the design of an all-purpose model which can adapt depending on the environment requires dealing with the polarizability issue, if not exactly, at least approximately. Unfortunately, polarizable models such as the BK3 Kiss and Baranyai (2013), HBP Jiang et al. (2016), and iAMOEBA Wang et al. (2013) give only a modest improvement, but their computational cost makes them inefficient for large system sizes. On the other hand, the very active line of research based on machine learned potentials can do no better than the underlying approximations Zhang et al. (2018), and some recent studies for state of the art density functionals provide very disappointing results Montero de Hijes et al. (2023, 2024). Further improvement along this line requires a training set obtained from expensive ab initio calculations dealing accurately with electronic correlations, but alas, neural networks appear to exhibit here significant limitations in modelling the ab initio potential energy surface for a wide range of different chemical environments Zhai et al. (2023).
@@
Due to these difficulties, model development is still often directed to the design of simple and tractable force fields, based on intuition and basic principles of chemistry. As a result, one often neglects polarizability altogether and describes the average molecular polarity with a simple distribution of fixed point charges embedded into a spherical core. This is the case of the celebrated SPC model Berendsen, Grigera, and Straatsma (1987). Here, two positive partial charges sitting on hydrogen positions are neutralized by a negative charge located at the center of a simple Lennard-Jones site, which accounts for the repulsion and dispersion of the oxygen atom. Alternatively, the negative charge can be shifted along the bisector axis of the HOH angle, as proposed by Bernal-Fowler already in 1933 Bernal and Fowler (1933), leading to the so-called TIP4P family of water models.Jorgensen et al. (1983)
@@
Despite their simplicity, SPC and TIP4P models do a rather good job at describing the qualitative behavior of liquid water at ambient conditions, but the TIP4P family has been shown to be more robust and remains qualitatively correct over a much wider range of thermodynamic conditions Sanz et al. (2004); Abascal et al. (2007). Particularly, the TIP4P/2005 model provides a good qualitative description of the phase diagram, including the coexistence between several solid phases, as well as a quantitative description of the temperature of maximum density, the orthobaric liquid densities, molar volumes and surface tensions Abascal et al. (2005, 2007).
@@
Of course, neglecting polarizability comes at a price. For the particular case of the TIP4P/2005 model, the accuracy of liquid phase densities is at the cost of a rather poor description of the dielectric constant, which is about 30% too low compared to experiments Vega and Abascal (2011). Alternatively, it often occurs that models providing a good description of the dielectric constant, such as SPC/E or TIP3P Berendsen, Grigera, and Straatsma (1987); Jorgensen et al. (1983), yield a qualitatively incorrect description of the phase diagram Vega and Abascal (2011).
@@
A simple way out of this problem was suggested by Leontyev and Stuchebrukhov some time ago Leontyev and Stuchebrukhov (2009, 2011). These authors argued that the electric field generated by point charges on a molecule is screened by the electronic polarizability of the surrounding molecules. Therefore, the "effective" charges, that are required to reproduce an accurate force field are smaller than the true point charges of the molecule, by a factor equal to the high frequency refractive index, , i.e. . Consequently, models predicting a good force field will yield too low a molecular dipole moment when measured in terms of the effective charge. However, assuming the true charge is a factor larger than the effective charge, and noting water’s refractive index is ca. , brings the dipole moment of usual force fields ( D) in close agreement with results of ab initio calculations ( D).
@@
This appealing idea has received great attention and has become known as the scaled charge paradigm Kann and Skinner (2014); Kirby and Jungwirth (2019); Jorge and Lue (2019); Zeron, Abascal, and Vega (2019); Han, Isborn, and Shi (2023); Weldon and Wang (2023); Blazquez, Conde, and Vega (2023). As another side of the same coin, some authors have stressed the need to make a distinction between the potential energy surface, as governed by , and the so called dipole moment surface, which is dictated by Vega (2015). This concept can be useful to bring predicted dielectric constants into agreement with experiment Aragones, MacDowell, and Vega (2011a); Han, Isborn, and Shi (2023), but must be exercised with great caution. Indeed, these two surfaces could be different but are definitively not independent. Linear response theory implies that a molecule cannot tell the difference between an external electric field (as dictated by a test charge) or that emanating from a neighboring molecule within the system (as dictated by the effective charge). However, in the scaled charge paradigm the dipolar response would be dictated by the true dipole in the former case and by the scaled dipole in the latter. As a result, arbitrarily scaling the molecular dipole to bring the dielectric constant into agreement with experiments is likely to violate the dissipation-fluctuation theorem.
@@
Be as it may, the charge scaling concept serves to increase the number of liquid state properties that can be predicted by non-polarizable force fields, and optimized models such as TIP4P/2005 can be considered as very robust tools for the prediction of pure water properties. However, the lack of polarization will result in poor performance in cases where environments with widely differing structures are encountered in the same system Abascal et al. (2007).
@@
One such important situation is when two or more phases need to be considered simultaneously, as in phase coexistence. Indeed, it has been widely documented that it is not possible to predict accurate liquid water properties while providing a good estimate of the melting point Abascal et al. (2007). Similarly, point charge models provide usually rather poor vapor pressures, even when including ad hoc corrections for the free energy of polarization Berendsen, Grigera, and Straatsma (1987); Vega, Abascal, and Nezbeda (2006).
@@
A second very significant case involves electrolyte solutions, where ions can create very strong local electric fields and polarization becomes then a major issue. Indeed, for a single isolated ion of charge , the free energy of solvation is proportional to Israelachvili (1991), so one can anticipate that the lack of an accurate dielectric constant will result in a poor description of the solution. Not surprisingly, attempts to reproduce properties of electrolyte solutions using the expected nominal charge have met only limited success Joung and Cheatham (2008); Yagasaki, Matsumoto, and Tanaka (2020); Smith and Dang (1994); Fuentes-Azcatl and Barbosa (2016).
@@
The scaled charge concept comes here at the rescue, under the assumption that also the nominal charge of the ions needs to be scaled Kann and Skinner (2014). Indeed, several authors have recently exploited this idea Kann and Skinner (2014); Kirby and Jungwirth (2019); Jorge and Lue (2019); Zeron, Abascal, and Vega (2019); Han, Isborn, and Shi (2023); Blazquez, Sedano, and Vega (2024), and showed that some of the TIP4P models with poor predictions of the dielectric constant can provide rather accurate predictions of a number of solution properties by scaling the nominal charges by a factor of ca. Zeron, Abascal, and Vega (2019); Blazquez et al. (2022); Trejos et al. (2023); Blazquez, Conde, and Vega (2023). Unfortunately, predictions here are far less robust than for pure water, and different scaling factors need to be considered depending on whether one seeks for accurate predictions of bulk thermodynamic properties, surface tensions, or transport properties Blazquez, Conde, and Vega (2023); Le Breton and Joly (2020). Moreover, the improvement of the solution properties comes at the cost of significantly deteriorating the description of the melt and the crystal, which also results in poor predictions of the electrolyte’s solubility.
@@
For this reason, some authors have sought for an improved description of solution properties by using water’s dielectric constant as a major target property in the model parametrization Izadi, Anandakrishnan, and Onufriev (2014); Fuentes-Azcatl and Alejandre (2014)
@@
This is the case of the OPC modelIzadi, Anandakrishnan, and Onufriev (2014), which has been recently recommended for use with the Amber package Tian et al. (2020). Unfortunately, the improved description of the dielectric constant in this model is at the cost of using the location of the hydrogen sites as fitting parameters. This can provide extra flexibility for the parametrization, but will obviously ruin the molecule’s moment of inertia, which is one of the very few molecular properties that we can safely constrain from the outset. Of course, the dynamics will suffer from such a choice, and recent work has shown that the increased parameter space chosen is not sufficient to improve other models such as TIP4P/2005 Sedano, Blazquez, and Vega (2024). On the other hand, TIP4P/, a model developed by Fuentes-Azcatl et al. at about the same time Fuentes-Azcatl and Alejandre (2014), reproduces experimental dielectric constants over a wide range of thermodynamic conditions, with only small variations of the TIP4P/2005 charges and geometry of the parent model. As a result, it can also predict accurately the location of the temperature of maximum density. The simultaneous prediction of these two properties is rather uncommon in point charge force fields and appears to suggest a good performance of solution properties.
@@
Based on the scaled charge paradigm, however, it would appear that using dielectric constants as a target property is unsafe and will result in a poor description of the"true" force field Sedano, Blazquez, and Vega (2024). According to this argument, one would then expect that the improved performance will be at the cost of spoiling other properties, as it appears to be the case for the OPC model Sedano, Blazquez, and Vega (2024).
@@
In this work, we assess the performance of the TIP4P/ water model for a wide range of different properties and compare its performance with OPC and TIP4P/2005 using the well known benchmark proposed by Abascal and Vega (VA) Vega and Abascal (2011). The model is then tested for transferability by studying solution properties as described by the Joung-Cheatham electrolyte force field Joung and Cheatham (2008), which was not specifically parametrized for TIP4P/. Our results show that TIP4P/ performs essentially as well as TIP4P/2005 for pure water and that it remains a transferable and robust model for the study of electrolyte solutions without additional parametrization. This shows that a robust parametrization can be achieved without invoking the appealing concept of charge scaling.
II Overview of charge scaling
@@
Such a continuum description makes sense provided the distance between the interacting charges is large compared with the typical distances between the polarizable solvent molecules. Charge scaling is therefore the exact long wavelength result for the interaction between ions in a solvent. Assuming that interactions between nearby ions or solvent molecules are mediated by a dielectric continuum is likely to be significantly less accurate, however.
@@
Assuming an isotropic, local and linear response of the dielectric background, the polarization is , where is the background’s susceptibility. In as much as the nuclei are fixed, and the polarization is only due to the electronic degrees of freedom, we can assume is dictated by the static electronic polarization of the system. In practice, electronic degrees of freedom are unresponsive below an optical frequency, Parsegian (2005); Israelachvili (1991). Therefore, the response of the medium is dictated by the system’s index of refraction in the visible, , which is related to the high frequency dielectric constant as .
@@
Notice that we have taken care to write here the screening felt by charges in terms of the solvent’s refraction index in order to make explicit that it is only the electronic degrees of freedom that participate, i.e., the screening here does not account for the low frequency permanent dipole fluctuations of the solvent, and it is more akin to than to the static dielectric constant .
Eq.(2) provides the rationale for charge scaling. Since the pair energy is dictated by the scaled charges, , and the dielectric constant measures fluctuations of the actual charges, , it would appear to be unwise to use dielectric constants as a target property in force field modeling.
@@
where the non-local electronic response is now given in terms of the wave-vector dependent susceptibility as .
This shows implicitly that the extent of screening felt by the molecules is position dependent. In order to illustrate this explicitly, a model for the complex refraction index, is needed. For simple models of water and dipolar hard spheres,Stell, Patey, and Høye (1981); Trokhymchuk, Holovko, and Heinzinger (1993); Bopp, Kornyshev, and Sutmann (1996, 1998); Seyedi, Martin, and Matyushov (2019); Becker et al. (2025) simulations and integral equation theory show that the wave-vector dependent dielectric constant has a rather complex behavior, and exhibits at least two poles at finite wave-vectors. However, those calculations refer to the static dielectric constant, which is dominated by the fluctuations due to permanent dipole moments of the solvent.Stell, Patey, and Høye (1981) Here, we are only interested in the contribution of electronic fluctuations for frozen realizations of the nuclear positions, which is a problem far less well understood. In the absence of a suitable theoretical framework, we suggest here a minimal model, meant to illustrate how the distance dependent screening could emerge from the wave-vector dependent refractive index. For this purpose, we consider that two very nearby charges in a dilute gas effectively feel no screening, while two very distant charges will be screened as in a continuum. Assuming an even power dependence borrowed from statistical mechanics and field theory,Hansen and McDonald (1986); Trokhymchuk, Holovko, and Heinzinger (1993); Bopp, Kornyshev, and Sutmann (1996); Seyedi, Martin, and Matyushov (2019) a simple interpolating formula between these two limits is:
where is a correlation length in the order of the solvent molecular spacing. This equation resembles the Random Phase Approximation for the one component plasma, but with a cutoff to avoid the low wave-vector divergence that is typical in metals.Hansen and McDonald (1986)
Fourier transforming Eq.(4), using Eq.(5) for the polarization response yields an effective pair potential for the electronic background mediated interactions (cf. Mondal and Bagchi (2024)):
This result smoothly interpolates between Coulomb’s law in a vacuum and a screened Coulomb’s law for charges in a polarizable background. The crossover takes place at a microscopic length scale of . This result might be a mere caricature of the complex distance dependent screening in water Bopp, Kornyshev, and Sutmann (1996, 1998); Berthoumieux (2018), but helps to show how the simplest account of non-locality can lead to deviations from the continuum approximation. Interestingly, Eq.(6) can be also interpreted as describing a plain Coulomb law with an dependent permittivity of , which smoothly switches from unscreened interactions to the expected long wave-length limit for large . The crucial issue here is what is the extent of charge screening at the typical distance of the closest approach between the molecules, say, . i.e., whether is already sufficiently small for full screening to have set on. Since we expect on the scale of , this is not likely to be the general situation. This is in line with recent ab initio calculations between ions dissolved in argon, which show that the screening at contact distances is not quite that achieved at large distances, but at least, definitively much larger than that expected for ions in vacuum Kostal, Jungwirth, and Martinez-Seara (2023). This means that charge screening is not fully warranted at contact distances, but could be a better approximation than assuming charges interacting in vacuum.
III Methods
III.1 Dielectric constant